Q:-Let the points A9a,b,c) and B(d,e,f) be at distances p and q from origin. The line AB passes through origin when


A. (d/a)=(e/b)=(f/c)


B. ad+be+cf=pq


C. ad+be+cf=p2+q2


D. None of these

Grade:12

1 Answers

Badiuddin askIITians.ismu Expert
14 years ago

Dear Tapasranjan

AB line passes through origin so slope of line joining the origin and A  , and slope of line joining the origin and B will be same

d-0 /a-0   = e-0 /b-0   = f-0 /c-0

or

 d/a = e/b  =f/c = k   .....................1

option a is correct

now use distance formula

a2 +b2+ c2 = p2   

and

d2 + e2 + f2 =q2

now

 put value of d,e,f in terms of a,b,c

 k2 ( a2 +b2+ c2 ) = q2 

 k2 ( p2 ) = q2  

 k = q/p


now let  ad +eb + fc

          = k(a2 +b2+ c2 )

          =q/p  ( p2)

          =pq

  option b is also correct

Please feel free to post as many doubts on our discussion forum as you can.
If you find any question Difficult to understand - post it here and we will get you the answer and detailed
solution very  quickly.
 We are all IITians and here to help you in your IIT JEE  & AIEEE preparation.

 All the best.
 
Regards,
Askiitians Experts
Badiuddin